# Vision foundation models: can they be applied to astrophysics data?

### v3

@@ Abstract
Vision foundation models, which have demonstrated significant potential in multimedia applications, are often underutilized in the natural sciences. This is primarily due to mismatches between the nature of domain-specific scientific data and the typical training data used for foundation models, leading to distribution shift. Scientific images can have unique structure and characteristics, and researchers often face the challenge of optimizing model performance with only a few hundred or thousand labeled examples. Furthermore, the choice of vision foundation model can be non-trivial, as each model exhibits unique strengths and limitations, influenced by differences in architecture, training procedures, and the datasets used for training. In this work, we evaluate the application of various vision foundation models to images from optical and radio astronomy. Our results show that use of certain foundation models improves the classification accuracy of optical galaxy images compared to conventional supervised training. Similarly, we achieve equivalent or better performance in object detection tasks with radio images. However, their performance in classifying radio galaxy images is generally poor and often inferior to traditional supervised training results. These findings suggest that selecting suitable vision foundation models for astrophysics applications requires careful consideration of the model characteristics and alignment with the specific requirements of the downstream tasks.
@@ 1 Introduction
In recent years, foundation models have enabled significant advancements in both language and image processing. These large models, trained on vast amounts of data encompassing multiple domains, serve as a cornerstone for numerous everyday applications. Foundation models are designed either for representation learning, where the goal is to capture essential characteristics of the data, or for generative purposes, where the model attempts to produce new data samples similar to the training distribution. In this work, we focus on the potential of the learned representations from vision foundation models applied to astrophysical images.
A variety of architectural innovations and training methodologies have been employed to capture complex patterns in large datasets of natural images, some scraped from the internet and others, like ImageNet [12], more curated. For many computer vision applications, this provides a robust starting point for further adaptation via fine-tuning for specialized downstream tasks (DSTs). However, the diversity and complexity inherent in foundation models pose substantial hurdles for scientists in fields such as physics, astronomy, and biology, who wish to apply these tools for domain-specific inquiries.
@@
The objectives and data used for scientific DSTs are neither known nor used during the training of foundation models. Foundation models are typically trained in a self-supervised or weakly supervised way. Thus, in accordance to the information bottleneck principle [43], it is not obvious whether these models are able to retain features related to the DST, namely sufficient statistics, in their learned representations. If, as suggested, only DST-related information should be kept in the model’s representations (and the rest filtered out), it is difficult to determine what, if any, relevant information a foundation model trained on ImageNet might have for astronomy. Furthermore, since there is likely a misalignment between the training data distribution and the DST data distribution, this can result in the problem of distribution shift.
In addition to theoretical concerns, practitioners face the challenge of finding optimal fine-tuning techniques. Given that labeled data available for scientific applications is typically limited, identifying the appropriate training objective, data augmentations, model architecture, and training hyperparameters is essential to enhance the model’s performance. Firstly, there is the challenge of selecting an appropriate model, which is both critical and non-trivial, requiring a deep understanding of the model’s compatibility with the DST data distribution. Secondly, performance variation across different tasks and datasets can be perplexing. Some models excel with minimal labeled data, while others might require substantial label-rich datasets to perform effectively. This performance disparity is often not linear and depends on the nature and volume of the training examples provided for the DST. Thirdly, the feasibility of deploying these models is constrained by available computational capabilities, which varies dramatically based on the model’s scale —- from small to giant models. Finally, all of this investigation requires time and substantial effort from domain experts in both the topic of interest and in machine learning best practices.
The goal of our paper is to address this first challenge by investigating the differences between foundation models and the implications for obtaining optimal task-specific performance. Our experiments used images from both optical and radio astronomy, and we examined classification and detection tasks. By applying various foundation models to these datasets and DSTs, this paper seeks to bridge the gap between advanced machine learning techniques and practical scientific research.
@@ 2 Use of Foundation Models in Astrophysics
Figure 1: Random samples from each of the following datasets: ImageNet-1K (top row), GalaxyMNIST (second row), Radio Galaxy Zoo (RGZ, third row), MeerKAT MGCLS (bottom row). GalaxyMNIST and RGZ are labeled according to morphology class, while MGCLS is labeled by the number of compact sources present. White boxes indicate object bounding boxes used for source detection, when applicable. The rightmost two columns display sample images in their raw data format, unscaled, with an accompanying histogram. GalaxyMNIST combines data from the Dark Energy Camera’s r, g, and z channels, while radio images are single-channel continuum images reconstructed from Fourier space, which can therefore include negative values.
Figure 1 illustrates the differences between natural images that comprise common training datasets like ImageNet and images from optical and radio astronomy. Unlike natural images, astrophysics images tend to have the following properties:
Sparseness: most images consist of several objects that occupy a small fraction of the total image. Histograms of radio and optical images demonstrate how many pixels have very low brightness.
@@
High dynamic range: the brightness of the objects in the image can span several orders of magnitude, so traditional normalization is difficult. The ideal combination of filters or scaling for image display is often determined by research goals (e.g., emphasizing diffuse emission over compact sources).
Artefacts: instrumental effects or residuals from image reconstruction can form spurious structures of different scales in the images.
Because of these fundamental differences, it is commonly assumed in the academic community that vision models trained on natural images are inadequate for performing tasks in astronomy, especially radio astronomy, where images are mathematically reconstructed from sparse samples measured in the Fourier plane.
While there are many extremely specialized DSTs in both optical and radio astronomy, often taking advantage of the many wavelength channels available, some more common tasks involving a single or small number of channels are listed in Table 1. Machine learning can offer a number of applicable techniques, and pretrained foundation models can be of use in most cases.
Astrophysics Task Machine Learning Task or Method References
Image reconstruction CNNs, de-noising diffusion [39, 13]
Source detection Object detection [48, 20, 38]
Source characterization Object segmentation [15, 40]
Source classification Image or object classification [2, 38, 31]
Source deblending Instance segmentation [2, 36, 17]
Object/event discovery Anomaly detection [23, 45]
RFI detection CNNs, GANs [46, 27]
Background/foreground removal UNET, de-noising diffusion [6, 53, 5]
@@ 2.1 Data
In this work we evaluated galaxy morphology classification and source detection, two tasks common to both optical and radio astronomy. Details of the datasets used to perform these tasks are in Table 2, and sample images, along with histograms of their un-normalized original images, are in Figure 1.
Classification datasets are images with a single galaxy centered in the cutout. GalaxyMNIST (GMNIST, [51]) is a balanced dataset of four categories: smooth and round (SR), smooth and cigar-shaped (SC), edge-on-disk (E), and unbarred spiral (U). Radio Galaxy Zoo (RGZ) is unbalanced, with labels determined according to number of distinct radio components (C) and number of intensity peaks (P) in each source [49]. Possible combinations are: 1C 1P, 1C 2P, 1C 3P, 2C 2P, 2C 3P, and 3C 3P, a total of 6 classes. Labeling was done by citizen scientists who performed inspection by varying the relative intensity of continuum radio emission and infrared observations. It is not always visually obvious from the normalized PNG images if there is a difference between a cutouts containing the same number of bright peaks – for example, one containing a single galaxy with two peaks, and another two different components with one peak each. Image pre-processing was by necessity done differently than for natural images, with techniques such as below-threshold zero replacement or channel-specific normalization. Full details are given in their documentation, and for classification both datasets are used as-is, with no additional pre-processing performed.
@@
RGZ222[49] 7.7K $\times×132$ noisy, centered on bright radio galaxy classification, detection
@@
Source detection datasets also consist of cutouts from wide-field images; they may have one to six galaxies in the central area, as in RGZ, or have tens of galaxies in a single cutout, as in MGCLS. Examples of bounding boxes around sources are shown in Figure 1. MGCLS labels are consistent in that the boxes only designate compact sources, and not other examples of extended emission or larger sources that might also be present in the images. Because RGZ’s labels also contain extended sources, bounding boxes can much larger and filled with a large amount of noise.
@@ 2.2 Foundation Models
Foundation models exhibit significant variability in terms of their architectures, training, and performance characteristics. These differences arise from factors such as the size and nature of the training datasets, the number of parameters they incorporate, and their underlying architectural frameworks. Moreover, these models leverage a carefully curated set of data augmentation techniques and diverse pretraining strategies, each tailored to minimize a specific loss function. Table 3 lists the foundation models investigated in this study. With the exception of DINOv2, all were pretrained on ImageNet-1k’s 1.2 million training images.
@@
MAE [16] ViT-Base 16x16 86M ImageNet-1k -
DINOv2 [33] ViT-Base 14x14 86M LVD-142M (proprietary) +
DINOv1 [9] ViT-Base 16x16 86M ImageNet-1k +
MSN [1] ViT-Base 16x16 86M ImageNet-1k +
ResNet 50 [19] ResNet-50 25.6M ImageNet-1k -
ResNet 18 [19] ResNet-18 11.5M ImageNet-1k -
@@
MAE. The Masked AutoEncoder (MAE, [16]) uses masked image modeling (MIM) pretraining, reconstructing masked image patches from only a few visible patches. Unlike NLP inspired approaches like BeIT [3], which discretize visual tokens through an autoencoder, MAE directly processes the visible image patches through an encoder. The resulting output is then combined with mask tokens to reconstruct the original image using a decoder. Reconstruction error is used as the objective training function.
DINO. DINOv2 [33] improves upon its predecessor, DINOv1 [9], by utilizing a larger and more curated dataset, LVD-142M. DINOv2 integrates the DINO cross-entropy loss with the MIM objective employed in iBOT [54]. DINOv2 benefits from the Sinkhorn-Knopp batch normalization technique used in SwAV [7]. The model employs a teacher-student training paradigm, where a teacher network computes the feature representation of global views of an image, while a student network computes the feature representation of local views, a series of smaller crops. The model is optimized to train the student to replicate the teacher’s output, and the teacher is periodically updated using an exponential moving average (EMA) of the student’s parameters.
MSN. Masked Siamese Networks (MSN, [1]) combine MIM with Siamese networks to avoid pixel-level and token-level reconstructions. MSN also uses teacher-student training, with the student network computing the feature representation of a partially masked image.
ResNet. Residual Networks are a staple of computer vision, introduced by in [19]. They are convolutional neural networks (CNNs) that use skip connections to bypass one or more layers, allowing the network to maintain accuracy over very deep networks. ResNet18 consists of 8 blocks, while ResNet50 has 16. We use the pre-trained weights available through torchvision [28], which were obtained by training in a supervised fashion for classification of ImageNet-1k using cross-entropy loss. A series of very specific data augmentations, including TrivialAugment [29], random erasing, mixup and cutmix, helped increase top-1 accuracy relative to the original augmentation scheme.

### v4

# Bridging the Gap: Examining Vision Foundation Models for Optical and Radio Astronomy Applications

@@ Abstract
Context. Vision foundation models, which have demonstrated significant potential in many multimedia applications, are often underutilized in the natural sciences. This is primarily due to mismatches between the nature of domain-specific scientific data and the typical training data used for foundation models, leading to distribution shifts. Scientific data often differ substantially in structure and characteristics, and researchers frequently face the challenge of optimizing model performance with limited labeled data of only a few hundred or thousand images.
Aims. This work evaluates the performance of vision foundation models in astrophysics, with a focus on identifying the best practices for adapting these models to domain-specific datasets. We aim to establish a framework for selecting, fine-tuning, and optimizing these models for common tasks in optical and radio astronomy.
Methods. We compared multiple foundation models, including self-supervised, weakly supervised, and distillation-based architectures, across two representative optical and radio datasets. Experiments involved different fine-tuning strategies, projector heads, and data preprocessing techniques, with performance evaluated on classification and detection metrics.
Results. Features extracted by specific foundation models improved classification accuracy for optical galaxy images compared to conventional supervised training. Similarly, these models achieved equivalent or superior performance in object detection tasks with radio images. However, classification performance for radio galaxy images was generally poor, often falling short of traditional supervised approaches.
Conclusions. These findings suggest that selecting suitable vision foundation models for astrophysics applications requires careful consideration of the model characteristics and alignment with the specific requirements of the downstream tasks. This study demonstrates that vision foundation models can be effectively adapted to astrophysical applications, provided practitioners iterate on model selection, training strategies, and data handling. The proposed framework bridges the gap between these advanced models and the unique demands of astronomy, enabling broader adoption of deep learning in the field.
Key Words.:
Techniques: image processing, Methods: data analysis, Radio continuum: general
@@ 1 Introduction
Large machine learning models, trained on vast amounts of data encompassing multiple domains, have in recent years enabled significant advancements in both language and image processing. Known as foundation models, these serve as a cornerstone for numerous everyday applications. Foundation models are designed either for representation learning, in which the goal is to capture essential characteristics of the data, or for generative purposes, where the model attempts to create new data samples similar to the training distribution. In this work, we focus on the potential of the learned representations from vision foundation models applied to astrophysical images.
While machine learning is frequently used to manage large amounts of astrophysics image data, vision foundation models are not yet commonly employed. With the exception of ResNets (He et al. 2016) pre-trained on ImageNet (Deng et al. 2009), most works in optical and radio astronomy employ models trained from scratch (e.g. Slijepcevic et al. 2024; Lochner & Bassett 2021; Becker et al. 2021). This includes networks trained for tasks such as image classification and object detection that should be able to benefit directly from the many foundation models available. A more detailed literature review will follow in Sections 3 and 4, but two major themes emerge: firstly, models trained on natural images are not believed to have learned features relevant to astrophysics images, and so are disregarded, and secondly, the desire to employ lightweight models as to not over-fit small labeled datasets precludes the use of modern foundation models, whose smallest architectures tend to have tens of millions of trainable parameters.
Indeed, the proliferation of foundation models in the recent years has led to a seeming overabundance of choice to the machine learning practitioner. Each model has its strengths and weaknesses; the diversity and complexity inherent in foundation models can pose substantial hurdles for scientists who wish to apply them for domain-specific inquiries. Therefore, the primary goal of this paper is to experimentally address the initial choice of vision foundation model, given properties of astrophysics data, specific downstream tasks (DSTs), and availability of resources for fine-tuning the foundation model.
Vision foundation models fall broadly into three categories, distinguished by their training methods, architectures, and objectives: self-supervised models trained on a single modality, weakly supervised multimodal models, and distillation models that agglomerate learned representations from multiple pre-trained models. Supervised models are some of the classic pre-trained models available through software libraries such as huggingface, Tensorflow, and PyTorch; convolutional neural networks (CNNs) trained for image classification like the ResNet family, and more specialized architectures trained for object detection and segmentation like YOLO (Redmon et al. 2016) and SAM (Kirillov et al. 2023). Self-supervised foundation models do not have a single DST objective, and include Masked AutoEncoders (MAE, He et al. 2022), Masked Siamese Networks (MSN, Assran et al. 2022), DINO and its improvement DINOv2 (Caron et al. 2021; Oquab et al. 2024). Weakly supervised models enable cross-modal representation learning, first introduced by CLIP (Radford et al. 2021), with its loss function improved upon by SigLIP (Zhai et al. 2023). Knowledge distillation is frequently used to extract smaller models from larger ones, but it can also unify representations of multiple models. AM-RADIO (Ranzinger et al. 2024) distills a student model from multiple teacher models, allowing it to integrate information from several existing foundation models including CLIP and DINOv2.
@@
There are also many types of loss functions used to train foundation models, which impact their effectiveness for specialized DSTs. Self-supervised learning frequently uses a contrastive loss to distinguish between different views of the data, or a reconstruction objective to reconstruct parts of an image. Non-contrastive methods, often employing regularization, are aimed at learning stable, robust features without explicit contrastive pairs. Foundation models also employ various backbone architectures with different numbers of parameters, ranging from convolutional networks (CNNs) to Vision Transformers (ViTs, Dosovitskiy et al. (2021)). Of more concern to scientists is the lack of scientific images in pre-training data. Most foundation models are pre-trained on ImageNet or similar large collections of natural images scraped from the web. Certainly the objectives and specific data used for scientific DSTs are neither known nor used during the training of foundation models. Thus, in accordance with the information bottleneck principle (Tishby et al. 2000), it is not obvious whether these models are able to retain features related to the DST, namely sufficient statistics, in their learned representations. If, as this principle suggests, only DST-related information should be kept in the model’s representations (and the rest filtered out), it is difficult to determine what, if any, relevant information a foundation model trained on ImageNet might have for astronomy; this is one potential explanation for the aforementioned lack of application of pre-trained foundation models in the literature. Furthermore, there is likely a misalignment between the training data distribution and the DST data distribution; this problem is known in the field of computer science as distribution shift.
In addition to theoretical concerns, practitioners face the challenge of finding optimal fine-tuning techniques. Given that labeled data available for scientific applications is typically limited, identifying the appropriate training objective, data augmentations, model architecture, and training hyperparameters is essential to enhance the model’s performance. In this work, we provide a framework to guide the selection of foundation models for astrophysical imaging, tailored to specific characteristics of the target dataset, downstream task, and available resources for fine-tuning. Through this empirical approach, our study aims to clarify the strengths and weaknesses of vision foundation models, as well as to provide practical insights that can inform model selection and adaptation.
The paper is organized as follows. First we discuss the use of foundation models in astrophysics in Section 2, along with our choices of illustrative datasets, various foundation models, and downstream task methodology. Section 3 examines galaxy morphology classification, and several fine-tuning techniques one can use to improve performance on a challenging dataset. A second DST, object detection, is evaluated in Section 4. Based on these experiments, we generalize our findings into practical guidelines in Section 5, before concluding in Section 6.
@@ 2 Use of Foundation Models in Astrophysics
Figure 1: Random samples from each of the following datasets: ImageNet-1K (top row), GalaxyMNIST (second row), Radio Galaxy Zoo (RGZ, third row), MeerKAT MGCLS (bottom row). GalaxyMNIST and RGZ are labeled according to morphology class, while MGCLS is labeled by the number of compact sources present. White boxes indicate object bounding boxes used for source detection, when applicable. The rightmost two columns display sample images in their raw data format, unscaled, next to the corresponding histogram. GalaxyMNIST combines data from the Dark Energy Camera’s r, g, and z channels, while radio images are single-channel continuum images reconstructed from Fourier space, which can therefore include negative values as indidcated by the different colors in the histogram.
Figure 1 illustrates the differences between natural images that comprise common training datasets like ImageNet and images from optical and radio astronomy. Unlike natural images, astrophysics images are found nowhere on Earth and tend to have the following properties:
Sparseness: most of the images consist of several objects that occupy a small fraction of the total image size.
@@
High dynamic range: the brightness of the objects in the image can span several orders of magnitude, not easily captured by traditional normalization methods. Research goals most often define the ideal combination of filters or various weighting schemes for displaying the image (for example, to emphasize compact bright sources vs diffuse extended emission).
Artefacts: instrumental effects or residuals from image reconstruction can form structures of different scales in the images.
Because of these fundamental differences, it is rare to encounter any deep learning applications in astrophysics literature where networks are trained starting from the pre-trained weights of a foundation model (the only example we found was Burke et al. (2019)). This is especially true for radio astronomy where images are mathematically reconstructed from sparse samples measured in the Fourier plane.
While there are many extremely specialized DSTs in both optical and radio astronomy, often taking advantage of the many wavelength channels available, some more common tasks involving a single or small number of channels are listed in Table 1. Machine learning can offer a number of applicable techniques, and it is possible to use pre-trained foundation models in most cases, despite the lack of examples in the literature.
Astrophysics Task Machine Learning Task or Method Selected References
Image reconstruction CNNs, de-noising diffusion (Schmidt et al. 2022; Drozdova et al. 2023)
Source detection Object detection (Vafaei Sadr et al. 2019; Jia et al. 2023; Riggi et al. 2023)
Source characterization Object segmentation (Farias et al. 2020; Sortino et al. 2023)
Source classification Image or object classification (Burke et al. 2019; Riggi et al. 2023; Merz et al. 2023)
Source deblending Instance segmentation (Burke et al. 2019; Reiman & Göhre 2019; Hausen & Robertson 2022)
Object/event discovery Anomaly detection (Lochner & Bassett 2021; Villar et al. 2021)
RFI detection CNNs, GANs (Vos et al. 2019; Li et al. 2021)
Background/foreground removal UNET, de-noising diffusion (Cohen & Lu 2021; Zhou et al. 2023; Chen et al. 2024a)
@@ 2.1 Data
In this work we chose to evaluate galaxy morphology classification and source detection, two tasks common to astronomy across the electromagnetic spectrum. Details of the datasets used to perform these tasks are in Table 2, and sample images are shown in Figure 1.
@@
RGZ222(Wong et al. 2024) 7.7K $\times×132$ noisy, centered on bright radio galaxy classification, detection
@@
Classification datasets are images with a single galaxy centered in the cutout. GalaxyMNIST (GMNIST, (Walmsley et al. 2022)) is a balanced dataset of four categories: smooth and round (SR), smooth and cigar-shaped (SC), edge-on-disk (E), and unbarred spiral (U). Radio Galaxy Zoo (RGZ) is unbalanced, with labels determined according to number of distinct radio components (C) and number of intensity peaks (P) in each source (Wong et al. 2024). The observed combinations present in RGZ are: 1C 1P, 1C 2P, 1C 3P, 2C 2P, 2C 3P, and 3C 3P, a total of 6 classes. Labeling was done by citizen scientists who performed inspection by varying the relative intensity of continuum radio emission and infrared observations. It is not always visually obvious from the normalized PNG images if there is a difference between a cutout containing the same number of bright peaks – for example, one containing a single galaxy with two peaks, and another two different components with one peak each.
Source detection datasets also consist of cutouts from wide-field images; they may have one to six galaxies in the central area, as in RGZ, or have tens of galaxies in a single cutout, as in MGCLS. Examples of bounding boxes around sources are shown in Figure 1; MGCLS labels are consistent in that the boxes only designate compact sources, and not other examples of extended emission or larger sources that might also be present in the images. Because RGZ’s labels also contain extended sources, bounding boxes can much larger and filled with a large amount of noise.
@@ 2.2 Foundation Models
Foundation models exhibit significant variability in terms of their architectures, training, and performance characteristics. These differences arise from factors such as the size and nature of the training datasets, the number of parameters they incorporate, and their underlying architectural frameworks. Moreover, these models leverage a carefully curated set of data augmentation techniques and diverse pre-training strategies, each tailored to minimize a specific loss function. Table 3 lists the foundation models investigated in this study. Most were pre-trained on ImageNet-1k’s 1.2 million training images.
Some foundation models are pre-trained with a particular downstream task objective, but even if they are not, they can be used as backbones together with particular projector heads that enable them to be leveraged for a variety of tasks. This involves removing any specialized head that is present for pre-training – for example, removing the final linear classifier layer of PyTorch’s pre-trained ResNet – and attaching the task-specific head.
2.2.1 Categories of foundation models
Earlier, we broadly categorized vision foundation models into three groups: self-supervised models (SSL), weakly supervised models (WSL), and distilled models. Each model type is characterized by different methodologies and architectures aimed at robust representation learning.
SSL models are designed to extract meaningful representations from large datasets without the need for labels, operating under the assumption that neither the downstream tasks nor the corresponding labels are available during training. This allows SSL models to focus on learning representations that are generalizable and transferable to many different tasks. Embeddings can be learned through masking regions or patches of the input image, then inferring the missing content as is done by Masked Siamese Networks (MSN). Embedding reconstruction models, such as Denoising Autoencoders and Masked Autoencoders (MAE), utilize an encoder-decoder architecture to compress the input image into a latent representation and reconstruct it. Alternatively, joint embedding models bypass reconstruction by learning representations for augmented views of the same image, with one encoder fixed as a reference and another trainable encoder processing a masked version. However, these models are susceptible to mode collapse in which embeddings lose meaningful variability. Techniques used by models like SimCLR (Chen et al. 2020), BYOL (Grill et al. 2020), MSN, SwAV (Caron et al. 2020), DINO, Barlow Twins (Zbontar et al. 2021), and VICReg (Bardes et al. 2022) can mitigate this issue. Hybrid models combining the principles of embedding reconstruction and joint embedding also exist, such as CAE (Chen et al. 2024b) and BeIT (Bao et al. 2022), and a recent innovation is joint embedding prediction models like I-JEPA (Assran et al. 2023) and World Models (Ha & Schmidhuber 2018), which introduce an additional projector to predict masked representations to enhance training stability.
Weakly supervised learning (WSL) models operate in scenarios where full labels are unavailable but auxiliary data, such as text descriptions or metadata, provides partial supervision. These models have dual-encoder architectures, most often with one processing images and another encoding text to generate semantic embeddings. The objective is to align these embeddings in a shared latent space. Representative models in this category include CLIP, GLIP (Li et al. 2022a), SigLIP, CoCa (Yu et al. 2022), and ALIGN (Jia et al. 2021), which excel in tasks like zero-shot learning and cross-modal retrieval.
Distilled models rely on a teacher-student paradigm, where a smaller, more efficient model (student) learns to replicate the outputs or distributions generated by a larger, more complex model or models (teacher). The goal is to transfer the teachers’ knowledge to the student model; distillation is most often used to create smaller versions of a large foundation model. It can also be used to combine and enhance the outputs of multiple teacher models, such that the student inherits their capabilities across multiple tasks. NVIDIA’s AM-RADIO model (Ranzinger et al. 2024) combines multi-teacher distillation with foundation models, distilling representations from various high-performance models in both standard vision and multi-modal tasks.
2.2.2 Foundation models used in this study
@@
MAE (He et al. 2022) ViT-Base 16x16 86M ImageNet-1k -
DINOv2 (Oquab et al. 2024) ViT-Base 14x14 86M LVD-142M (proprietary) +
DINOv1 (Caron et al. 2021) ViT-Base 16x16 86M ImageNet-1k +
MSN (Assran et al. 2022) ViT-Base 16x16 86M ImageNet-1k +
SigLIP (Zhai et al. 2023) ViT-Base 16x16 86M WebLI (English) -
AM-RADIO (Ranzinger et al. 2024) ViT-Base 16x16 98.2M various -
ResNet 50 (He et al. 2016) ResNet-50 25.6M ImageNet-1k -
@@
MAE. The Masked AutoEncoder (MAE) uses masked image modeling (MIM) pretraining, reconstructing masked image patches from only a few visible patches. Unlike natural language processing-inspired approaches like BeIT, which discretize visual tokens through an autoencoder, MAE directly processes the visible image patches through an encoder. The resulting output is then combined with mask tokens to reconstruct the original image using a decoder. Reconstruction error is used as the objective training function.
MSN. Masked Siamese Networks (MSN) combine MIM with Siamese networks to avoid pixel-level and token-level reconstructions. MSN also uses teacher-student training, with the student network computing the feature representation of a partially masked image.
DINO. DINOv2 improves upon its predecessor, DINOv1 (Caron et al. 2021), by utilizing a larger and more curated dataset, designated LVD-142M. DINOv2 integrates the DINO cross-entropy loss with the MIM objective employed in iBOT Zhou et al. (2022). DINOv2 benefits from the Sinkhorn-Knopp batch normalization technique used in SwAV. The model employs a teacher-student training paradigm, where a teacher network computes the feature representation of global views of an image, while a student network computes the feature representation of local views, a series of smaller crops. The model is optimized to train the student network to replicate the teacher network’s output. The teacher network is periodically updated using an exponential moving average (EMA) of the student network’s parameters.
SigLIP. Sigmoid language image pre-training (SigLIP), is a multimodal method designed to align image and textual representations in a shared embedding space using a modified contrastive learning framework. Pre-trained on WebLI’s English text-image pairs, SigLIP uses a sigmoid loss that allows the use of extremely large batch sizes in pre-training, when using multiple GPUs.
AM-RADIO. Agglomerative Model – Reduce All Domains Into One (AM-RADIO) merges CLIP, SAM, and DINOv2 into a unified model via multi-teacher distillation. Merging the concepts of a student learning from an ensemble of teachers with foundation models, AM-RADIO trained student models from CLIP, DINOv2, and SAM, the authors used a cosine distance loss to match a simple student adaptor head with the teacher spatial feature vectors and summary vectors (when available). In this way, the student is able to mimic the teacher and perform their targeted downstream tasks.
In addition to these models, we compare with ResNet, one of the most used networks in machine learning for astrophysics. Residual Networks are a staple of computer vision, introduced by He et al. (2016). They are convolutional neural networks (CNNs) that use skip connections to bypass one or more layers, allowing the network to maintain accuracy over very deep networks. ResNet-18 consists of 8 blocks, while ResNet-50 has 16. In this work we use the pre-trained weights available through torchvision (Marcel & Rodriguez 2010), which were obtained by training in a supervised fashion for classification of ImageNet-1k using cross-entropy loss. A series of very specific data augmentations, including TrivialAugment (Müller & Hutter 2021), random erasing, mixup and cutmix, helped increase top-1 accuracy relative to the original augmentation scheme of random resized crops and horizontal flips.